# Frontiers in Probabilistic Inference: Sampling Meets Learning

### v1

@@ D.2 Generalised self-consistency property
Define $\tilde{\mathbb{P}}$ to be the path measure associated to the auxiliary process ${\mathrm{d}}\tilde{X}_{t}=(b+\tilde{b})(t,\tilde{X}_{t}){\mathrm{d}}t+\sigma(t,\tilde{X}_{t}){\mathrm{d}}B_{t}$. For ease of notation, we will present the result for scalar constant diffusion coefficient $\sigma$; the proof of the general case follows with the appropriate changes. By Girsanov’s theorem [31, Theorem 8.6.8], the Radon-Nikodym derivative is
$\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}(X_{[s,T]})=\exp\bigg(-\int_{s}^{T}\sigma^{-1}\tilde{b}(r,X_{r})^{\top}{\mathrm{d}}B_{r}-\int_{s}^{T}\tfrac{\sigma^{-2}}{2}\|\tilde{b}(r,X_{r})\|^{2}\,{\mathrm{d}}r\bigg).$ (71)
As in the proof of Theorem 2.1, we begin with Doob’s $h$-transform, then perform the following calculations (where $B^{\mathbb{P}},B^{\tilde{\mathbb{P}}},B^{\mathbb{Q}}$ denote standard Brownian motions under their respective measures).
$u^{*}($ $s,x)=\nabla_{x}\log{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]$ (72)
$=\frac{1}{{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]}\nabla_{x}{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]$ (73)
$=\frac{1}{{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]}\nabla_{x}{\mathbb{E}}_{\tilde{\mathbb{P}}}[F(\tilde{X}_{T})\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}|\tilde{X}_{s}=x]$ (74)
@@
$\qquad+F(\tilde{X}_{T})\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}\Big(-\int_{s}^{T}\tfrac{1}{\sigma^{2}}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,\tilde{X}_{r})^{\top}\tilde{b}(r,\tilde{X}_{r}){\mathrm{d}}r\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,\tilde{X}_{r})^{\top}{\mathrm{d}}B^{\tilde{\mathbb{P}}}_{r}\Big)|\tilde{X}_{s}=x\Big]$ (75)
@@
$\qquad+F(X_{T})\Big(-\int_{s}^{T}\tfrac{1}{\sigma^{2}}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}\tilde{b}(r,X_{r}){\mathrm{d}}r$
$\qquad\qquad-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}({\mathrm{d}}B^{\mathbb{P}}_{r}-\tfrac{1}{\sigma}\tilde{b})\Big)|X_{s}=x\Big]$ (76)
$=\frac{1}{{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]}{\mathbb{E}}_{\mathbb{P}}\Big[\tilde{J}_{T|s}^{\top}\,\nabla F(X_{T})+F(X_{T})\Big(-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}{\mathrm{d}}B^{\mathbb{P}}_{r}\Big)|X_{s}=x\Big]$ (77)
$={\mathbb{E}}_{\mathbb{Q}}\Big[\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}({\mathrm{d}}B_{r}^{\mathbb{Q}}+\sigma u^{*}{\mathrm{d}}t)|X_{s}=x\Big]$ (78)
$={\mathbb{E}}_{{\mathbb{Q}}}\Big[\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})-\int_{s}^{T}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r|X_{s}=x\Big].$ (79)
Much of this derivation follows similarly to the standard case. When we perform the change of measures, we include the appropriate modifications to the Brownian motions; namely, we have ${\mathrm{d}}B^{\mathbb{P}}_{r}={\mathrm{d}}B^{\tilde{\mathbb{P}}}_{r}+\tfrac{1}{\sigma}\tilde{b}(r,X_{r}){\mathrm{d}}t$, and ${\mathrm{d}}B^{\mathbb{P}}_{r}={\mathrm{d}}B_{r}^{\mathbb{Q}}+\sigma u^{*}{\mathrm{d}}t$.
The remainder of the proof follows similarly to the proof of Theorem 2.1; as before, we ‘invert’ the property at a later time $t$, utilising the tower rule and the semigroup property of the Jacobian $\tilde{J}$.
$u^{*}($ $s,x)={\mathbb{E}}_{\mathbb{Q}}[\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})-\int_{s}^{T}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r|X_{s}=x]$ (80)
@@
$\qquad-\int_{t}^{T}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})|X_{t}=x]$ (81)
$\qquad-\int_{t}^{T}(J_{r|t}J_{t|s})^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+(J_{T|t}J_{t|s})\top\,\nabla\log F(X_{T})|X_{t}=x]$ (82)
@@
$\qquad-\int_{t}^{T}(\tilde{J}_{r|t}\tilde{J}_{t|s})^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+(\tilde{J}_{T|t}\tilde{J}_{t|s})^{\top}\,\nabla\log F(X_{T})|X_{[s,t]}]~|X_{s}=x\big]$ (83)
@@
$\qquad+\tilde{J}_{t|s}^{\top}\,{\mathbb{E}}_{\mathbb{Q}}[-\int_{t}^{T}\tilde{J}_{r|t}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+\tilde{J}_{T|t}^{\top}\,\nabla\log F(X_{T})|X_{t}]|X_{s}=x\big]$ (84)
$={\mathbb{E}}_{\mathbb{P}^{u^{*}}}\big[-\int_{s}^{t}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+\tilde{J}_{t|s}^{\top}\,u^{*}(t,X_{t})|X_{s}=x\big]$ (85)
To conclude, we show that the generalised self-consistency property retains the same theoretical results as in the standard case.
See B.2
Proof.
In light of the proof of Theorem D.2, we need only to show that the equivalence $(ii)\iff(iii)$ still holds. To do so, we mirror the proof of the standard case, and apply an Itô expansion of $\tilde{J}_{t|s}^{\top}\,u(t,X_{t}^{u})$ along the controlled dynamics,
${\mathrm{d}}(\tilde{J}_{t|s}^{\top}\,$ $u(t,X^{u}_{t}))=\tilde{J}_{t|s}^{\top}\,{\mathrm{d}}u(t,X^{u}_{t})+{\mathrm{d}}\tilde{J}_{t|s}^{\top}\,u(t,X^{u}_{t})$ (86)
$=\tilde{J}_{t|s}^{\top}\,\Big(\partial_{t}u(t,X_{t}^{u})\,{\mathrm{d}}t+\nabla u(t,X_{t}^{u})\cdot\left(b(t,X_{t}^{u})\,{\mathrm{d}}t+\sigma^{2}u(t,X_{t}^{u})\,{\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}\right)$
$\qquad+\tfrac{\sigma^{2}}{2}\Delta u(t,X_{t}^{u})\,{\mathrm{d}}t\Big)+\tilde{J}_{t|s}^{\top}\,\nabla(b+\tilde{b})(t,X_{t}^{u})\cdot u(t,X_{t}^{u}){\mathrm{d}}t$ (87)

### v2

@@ D.3 Generalisations of the Self-Consistency Property
In light of the proof of Theorem D.3, to prove Theorem 4.2 we need only to show that the equivalence $(ii)\iff(iii)$ still holds. To do so, we mirror the proof of the standard case, and perform an Itô expansion of $\tilde{J}_{t|s}^{\top}\,u(t,X_{t}^{u})$ along the controlled dynamics
${\mathrm{d}}(\tilde{J}$ $_{t|s}^{\top}\,u(t,X^{u}_{t}))=\tilde{J}_{t|s}^{\top}\,{\mathrm{d}}u(t,X^{u}_{t})+{\mathrm{d}}\tilde{J}_{t|s}^{\top}\,u(t,X^{u}_{t})$ (73a)
$=\tilde{J}_{t|s}^{\top}\,\left(\partial_{t}u(t,X_{t}^{u})\,{\mathrm{d}}t+\nabla u(t,X_{t}^{u})\cdot\left(b(t,X_{t}^{u})\,{\mathrm{d}}t+\sigma^{2}u(t,X_{t}^{u})\,{\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}\right)+\tfrac{\sigma^{2}}{2}\Delta u(t,X_{t}^{u})\,{\mathrm{d}}t\right)$ (73b)
$\qquad+\tilde{J}_{t|s}^{\top}\,\nabla(b+\tilde{b})(t,X_{t}^{u})\cdot u(t,X_{t}^{u}){\mathrm{d}}t$
[…]
The equivalence then follows by the same argument as in the proof of Theorem D.3, completing the proof of Theorem 4.2. The result of Theorem C.1 is also clear from (73c), by setting (Grad-HJB) to 0, integrating, and taking expectations. ∎
Connection to Reparameterisation Matrices
Consider a class of matrix-valued functions $M_{s}(t)$, with $M_{s}(t)$ invertible and $t\mapsto M_{s}(t)$ differentiable. By considering the above derivation but instead applying the Itô expansion to $M_{s}(t)u(t,X_{t}^{u})$, we see that the self-consistency property can in fact be generalised further to give
$M_{s}(s)u^{*}(s,x)={\mathbb{E}}\Big[\int_{s}^{t}\big(\partial M_{s}(r)-M_{s}(r)\nabla b\big)^{\top}\,u^{*}(r,X_{r}^{u^{*}}){\mathrm{d}}r+M_{s}(t)^{\top}\,u(t,X_{t}^{u^{*}})|X_{s}^{u^{*}}=x\Big].$ (74)
We focus in our presentation on taking $M_{s}(t)=\tilde{J}_{t|s}^{\top}$, as this provides a probabilistic interpretation of the role that these matrices play. Nevertheless, we provide this more general version here as it may be useful in some settings (and note that it justifies the use of (4.1) for spatially-dependent $\sigma$). The result relates to the family of ‘reparameterisation matrices’ introduced in Domingo-Enrich et al. (2024); we comment more on this connection in Appendix F.
Change-of-Measure Computation
For completeness, we here also fill in the details of the change-of-measure computation outlined in Section 4. This derives the generalised self-consistency property (40) in a way that mirrors the introductory derivation of (SC1) given in Section 2. Again, we present the computations for scalar $\sigma$ and the general result follows with the appropriate modifications. Note that as in the earlier case, this relation can also be obtained directly from (73c) in the proof above.
Define $\tilde{\mathbb{P}}$ to be the path measure associated to the auxiliary process ${\mathrm{d}}\tilde{X}_{t}=(b+\tilde{b})(t,\tilde{X}_{t}){\mathrm{d}}t+\sigma(t,\tilde{X}_{t}){\mathrm{d}}B_{t}$. By Girsanov’s theorem (Øksendal, 2003, Theorem 8.6.8), the Radon-Nikodym derivative is
$\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}(X_{[s,T]})=\exp\left(-\int_{s}^{T}\sigma^{-1}\tilde{b}(r,X_{r})^{\top}{\mathrm{d}}B_{r}-\int_{s}^{T}\tfrac{(\sigma\sigma^{\top})^{-1}}{2}\|\tilde{b}(r,X_{r})\|^{2}{\mathrm{d}}t\right).$
For ease of notation, we will present the result for scalar diffusion coefficient $\sigma$; the proof of the general case follows with the appropriate changes. As in the proof of Theorem 2.1, we begin with Doob’s $h$-transform, then perform the following calculations (where $B^{\mathbb{P}},B^{\tilde{\mathbb{P}}},B^{\mathbb{Q}}$ denote standard Brownian motions under their respective measures).
$u^{*}(s,x)$ $=$ $\nabla_{x}\log{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]$
$=$ $\frac{1}{{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]}\nabla_{x}{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]$
$=$ $\frac{1}{{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]}\nabla_{x}{\mathbb{E}}_{\tilde{\mathbb{P}}}[F(\tilde{X}_{T})\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}|\tilde{X}_{s}=x]$
@@
$\qquad\qquad+F(\tilde{X}_{T})\tfrac{{\mathrm{d}}\mathbb{P}}{{\mathrm{d}}\tilde{\mathbb{P}}}\Big(-\int_{s}^{T}\tfrac{1}{\sigma^{2}}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,\tilde{X}_{r})^{\top}\tilde{b}(r,\tilde{X}_{r}){\mathrm{d}}r-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,\tilde{X}_{r})^{\top}{\mathrm{d}}B^{\tilde{\mathbb{P}}}_{r}\Big)|\tilde{X}_{s}=x\Big]$
@@
$+F(X_{T})\Big(-\int_{s}^{T}\tfrac{1}{\sigma^{2}}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}\tilde{b}(r,X_{r}){\mathrm{d}}r-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}({\mathrm{d}}B^{\mathbb{P}}_{r}-\tfrac{1}{\sigma}\tilde{b}(r,X_{r}){\mathrm{d}}r)\Big)|X_{s}=x\Big]$
$=$ $\frac{1}{{\mathbb{E}}_{\mathbb{P}}[F(X_{T})|X_{s}=x]}{\mathbb{E}}_{\mathbb{P}}\Big[\tilde{J}_{T|s}^{\top}\,\nabla F(X_{T})+F(X_{T})\Big(-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}{\mathrm{d}}B^{\mathbb{P}}_{r}\Big)|X_{s}=x\Big]$
$=$ ${\mathbb{E}}_{\mathbb{Q}}\Big[\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})-\int_{s}^{T}\tfrac{1}{\sigma}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}({\mathrm{d}}B_{r}^{\mathbb{Q}}+\sigma u^{*}{\mathrm{d}}t)|X_{s}=x\Big]$
$=$ ${\mathbb{E}}_{{\mathbb{Q}}}\Big[\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})-\int_{s}^{T}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r|X_{s}=x\Big].$
Much of this derivation follows similarly to the standard case. When we perform the change of measures, we include the appropriate modifications to the Brownian motions; namely, we have ${\mathrm{d}}B^{\mathbb{P}}_{r}={\mathrm{d}}B^{\tilde{\mathbb{P}}}_{r}+\tfrac{1}{\sigma}\tilde{b}(r,X_{r}){\mathrm{d}}r$, and ${\mathrm{d}}B^{\mathbb{P}}_{r}={\mathrm{d}}B_{r}^{\mathbb{Q}}+\sigma u^{*}{\mathrm{d}}r$.
The remainder of the computation follows similarly to the proof of Theorem 2.1; as before, we ‘invert’ the property at a later time $t$, utilising the tower rule and the semigroup property of the Jacobian process $\tilde{J}_{t|s}$.
$u^{*}(s,x)$ $=$ ${\mathbb{E}}_{\mathbb{Q}}[\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})-\int_{s}^{T}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r|X_{s}=x]$
$=$ ${\mathbb{E}}_{\mathbb{Q}}[-\int_{s}^{t}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r-\int_{t}^{T}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+\tilde{J}_{T|s}^{\top}\,\nabla\log F(X_{T})|X_{t}=x]$
@@
$\qquad\qquad-\int_{t}^{T}(\tilde{J}_{r|t}\tilde{J}_{t|s})^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+(\tilde{J}_{T|t}\tilde{J}_{t|s})\top\,\nabla\log F(X_{T})|X_{t}=x]$
@@
$\qquad\qquad-\int_{t}^{T}(\tilde{J}_{r|t}\tilde{J}_{t|s})^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+(\tilde{J}_{T|t}\tilde{J}_{t|s})^{\top}\,\nabla\log F(X_{T})|X_{[s,t]}]~|X_{s}=x\big]$
@@
$\qquad\qquad+\tilde{J}_{t|s}^{\top}\,{\mathbb{E}}_{\mathbb{Q}}[-\int_{t}^{T}\tilde{J}_{r|t}^{\top}\,\nabla\tilde{b}(r,X_{r})^{\top}u^{*}(r,X_{r}){\mathrm{d}}r+\tilde{J}_{T|t}^{\top}\,\nabla\log F(X_{T})|X_{t}]|X_{s}=x\big]$
$=$ ${\mathbb{E}}\big[-\int_{s}^{t}\tilde{J}_{r|s}^{\top}\,\nabla\tilde{b}(r,X_{r}^{u^{*}})^{\top}u^{*}(r,X_{r}^{u^{*}}){\mathrm{d}}r+\tilde{J}_{t|s}^{\top}\,u^{*}(t,X_{t}^{u^{*}})|X_{s}^{u^{*}}=x\big],$
which completes the computation.
Appendix E Experimental Details and Additional Results
We implement our method using JAX (Bradbury et al., 2018). All experiments were carried out on a Nvidia GeForce RTX 2080Ti GPU unless otherwise specified. For the neural adjustment term $\eta_{\theta}$, we generally use a gradient-form parameterisation $\eta_{\theta}=\nabla\psi_{\theta}$ unless otherwise specified. We parameterise $\psi_{\theta}$ using the MLP architecture with sinusoidal time-embedding used in De Bortoli et al. (2021), with main hidden dimensions [128,128] and time-embedding hidden dimensions [64,64]. For fair comparison, we also used the same gradient-form architecture for NGDB, which we found performed slightly better than a standard architecture. When training NGDB, we used a Brownian auxiliary process. In all experiments, we use the Adam optimiser (Kingma & Ba, 2015). Below, we describe the individual experiment setups and implementation details used. For the alternative methods, we use the default hyperparameters provided in their respective codebases unless otherwise specified; we provide the links to these codebases in Appendix G.
Metrics
${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})={\mathbb{E}}\Big[\tfrac{1}{2}\int_{0}^{T}\lVert\sigma(t,X_{t}^{*})^{\top}u^{*}(t,X_{t}^{*})-\sigma(t,X_{t}^{*})^{\top}u_{\theta}(t,X_{t}^{*})\rVert^{2}{\mathrm{d}}t\Big].$ (75)
${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})={\mathbb{E}}\Big[\tfrac{1}{2}\int_{0}^{T}\lVert\sigma(t,X_{t}^{\theta})^{\top}u_{\theta}(t,X_{t}^{\theta})\rVert^{2}{\mathrm{d}}t\Big].$ (76)
In our experiments, we report values for these metrics computed using 1,000 trajectories. For the SDB method (Heng et al., 2025), we report these values for the forward bridge, which is learned by reversing the learned backwards bridge.
In all experiments, our method trained significantly faster than NGDB (see Table 2); this speed-up can be attributed to the fact that our method does not require backpropagation through the simulated trajectories. Both methods are implemented in JAX, allowing for fair comparison.
E.1 Ornstein-Uhlenbeck Process
${\mathrm{d}}X_{t}=-\alpha X_{t}{\mathrm{d}}t+\sigma{\mathrm{d}}B_{t},$ (77)
with starting state $x_{0}=e_{1}$, termination state $x_{1}=e_{2}$, termination time $T=1$, and diffusion coefficient $\sigma=0.1$. Ornstein-Uhlenbeck bridges are a standard setting in the literature for quantitatively evaluating performance, as the ground-truth control $u^{*}(t,x)$ can be calculated in closed-form as
$u^{*}(t,x)=\frac{2\alpha e^{-\alpha(1-t)}}{(1-e^{-2\alpha(1-t)})\sigma^{2}}(x_{1}-xe^{-\alpha(1-t)}).$ (78)
The training curves shown in Figure 3(a) are for $\alpha=2.0$ and $d=2$. The results in Figure 3(b) comparing the effect of dimension use $\alpha=2.0$, and the results comparing the effect of OU parameter in Figure 3(c) use $d=2$. To enable a fair comparison between the methods, we use the same gradient-form neural architecture for both, and use the same learning rate, number of gradient update steps, and number of trajectory simulations per step. We used 500 time-discretisation steps, simulated 64 trajectories at each step, and used a learning rate of 1e-4. For Figures 3(b) and 3(c), we ran the algorithms for 10,000 steps; we sometimes found the value for NGDB to start increasing again later in training, so the reported results are the lowest value for NGDB over the 10,000 training iterations.
Both methods learn the true drift to a very high degree of accuracy, and we remark that the difference in KL that is visible between the two methods is predominantly due to the behaviour towards the terminal time $T$. The methods behave comparably earlier in time, but the error of NGDB significantly increases approaching the endpoint. We hypothesise this is a consequence of the differing types of training objective; NGDB uses a KL-based objective, while our approach fits to the desired endpoint behaviour directly in the loss construction so is more accurate towards the terminal time. We were unable to learn the bridge dynamics successfully using the methods SDB, FB, and BEL, which is unsurprising given that the experiments involve rare events under the base dynamics.
E.2 Cox–Ingersoll–Ross model
The Cox–Ingersoll–Ross (CIR) model (Cox et al., 1985) is a 1-dimensional stochastic process that evolves according to the SDE
The bridges for this process are also known in closed form, as the transition densities can be written as $ce^{-u-v}\Big(\frac{v}{u}\Big)^{q/2}I_{q}(2\sqrt{uv}),$ where $c=\tfrac{2a}{(1-e^{-a(T-t)})\varepsilon^{2}},q=\tfrac{2ab}{\varepsilon^{2}}-1,u=cxe^{-a(T-t)},v=cx_{T}$, and $I_{q}$ is the modified Bessel function of the first kind of order $q$. This again allows us to report ${\mathrm{KL}}(\mathbb{P}^{*}\|\mathbb{P}^{\theta})$ and verify that the learned diffusion bridges are correct.
E.3 Double-Well Experiment
In the double-well experiments, we consider overdamped Langevin dynamics ${\mathrm{d}}X_{t}=-\nabla U(X_{t}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}$} using a classical double-well potential
$U(x)=v(x^{2}-1)^{2}.$
We aim to traverse from $x_{0}=1$ to $x_{1}=-1$ and set $\sigma=1.0$; this setting was previously considered in Pidstrigach et al. (2025). For each method, we used 500 time-discretisation steps and 4,000 training steps. For our method and NGDB we used a learning rate of 1e-3 and simulated 64 trajectories at each step, and we use the default hyperparameters for SDB and BEL. We note that we were unable to obtain good results for the double-well experiments using the FB method; we hypothesise that this is due to the behaviour of the adjoint process used for the training samples, which Baker et al. (2025) remark can exacerbate the difficulty of learning the correct dynamics for rare events.
The results reported in Table 1 are for a well height of $v=3.0$, which gives a strong peak between the wells. In Figure 5, we also compare the methods for different heights $v$ of the well barrier. We see that the methods that learn from controlled dynamics (that is, our method and NGDB) outperform SDB which learns from the unconditional dynamics; in particular, the performance of SDB becomes significantly worse for the largest value, which is unsurprising as the uncontrolled simulations will rarely cross between the wells in this case. For lower well heights, our method performs comparably with NGDB for both choices of self-consistency properties (SC1), (4.1). As the well height increases, we see that using (4.1) performs better; this reflects the discussion in Section 4 and demonstrates how avoiding large growth in the Jacobian terms can improve performance.
E.4 Cell Diffusion Process
Figure 6: Visualisation of the learned trajectories for each method in the normal, rare, and multimodal cell diffusion examples. The experimental setting is taken from Yang et al. (2025), using the cell differentiation model from Wang et al. (2011). Results in the leftmost column are obtained using the MCMC-based guided proposal approach from Schauer et al. (2017).
The cell differentiation model of Wang et al. (2011) has previously been used to provide a qualitative evaluation of diffusion bridge methods in Heng et al. (2025); Baker et al. (2025); Yang et al. (2025). The model dynamics are given by the 2-dimensional SDE
This provides an example in which the reference drift $b$ is not of gradient form, so following the discussion in Section 3 we include the reference drift in our parameterisation. We consider the three settings used in Yang et al. (2025): a normal event, a rare event, and a multi-modal example. For the normal event, the starting state is $x_{0}=[0.1,-0.1]$, the termination state is $x_{T}=[2.0,-0.1]$, and the termination time is $T=4$. For the rare event, the starting state is again $x_{0}=[0.1,-0.1]$, the termination state is $x_{T}=[1.0,-0.1]$, and the termination time is again $T=4$. For the multi-modal example, the starting state is $x_{0}=[-1.0,-1.0]$, the termination state is $x_{T}=[-1.0,-1.0]$, and the termination time is $T=5$. In all settings, the diffusion coefficient is $\sigma=0.1$, and we train for 1000 steps. For our approach and NGDB, we use a learning rate of 1e-3, and otherwise use the default parameters for the alternative methods.
As there is no ground-truth available in this setting, in Table 1 we report the KL divergence ${\mathrm{KL}}(\mathbb{P}^{\theta}\|\mathbb{P})$ relative to the base process. For a qualitative comparison, we additionally provide visualisations of the paths obtained by each method in the different cases in Figure 6. In the plot we also include a comparison with a non-ML baseline, displaying samples obtained using the MCMC-based guided proposal method from Schauer et al. (2017). We use the implementation and hyperparameters from Yang et al. (2025) (other than for the multimodal example, for which we used $\eta=0.995$ and ran several MCMC chains).
From the trajectories in Figure 6, we see that CCDB performs very similarly in each setting to NGDB, which was shown in Yang et al. (2025) to perform strongly in these experiments. Both methods accurately approximate the correct dynamics, even in the rare and multimodal examples. In contrast, the methods that learn from uncontrolled dynamics perform well in the normal setting, but struggle in the other cases as expected.
E.5 Müller-Brown Potential
We follow the experimental setup used in Du et al. (2024). The dynamics follow the overdamped Langevin SDE ${\mathrm{d}}X_{t}=-\nabla U(X_{t}){\mathrm{d}}t+\sigma{\mathrm{d}}B_{t}$}, where the potential $U$ is defined as
$U(x,y)=$ $-200\exp\big(-(x-1)^{2}-10y^{2}\big)$ (81)
$-100\exp\big(-x^{2}-10(y-0.5)^{2}\big)$ (82)
$+15\exp\big(0.7(1+x)^{2}+0.6(x+1)(y-1)+0.7(y-1)^{2}\big),$ (84)
with initial state $x_{0}=[-0.559,1.442]$, terminal state $x_{T}=[0.624,0.028]$, diffusion term $\sigma(t,x)=3\,{\mathrm{Id}}$}, and termination time $T=0.05$.
We used 1000 discretisation steps. For our method and for NGDB, we ran for 4000 training steps, simulated 64 trajectories at each outer step, and used a learning rate of 3e-4; for DL (Du et al., 2024) we use the default parameters provided in their codebase. For our method, we clipped the losses at 1e4, which we found to improve stability.
This experiment provides another example showcasing the benefit of using the self-consistency property (4.1) in our algorithm, compared to using the first version (SC1). This helps to avoid large growth in the Jacobian terms caused by the steep potential. Using (SC1) generally caused learning to be unstable and usually fail, other than for some specific choices of $\alpha$-schedule (see the trajectories in Figure 7). In contrast, using (4.1) gave consistent and reliable training.
(a) Using the self-consistency property (SC1) gives unstable learning due to the large Jacobian terms, and thus fails to learn the correct dynamics.
(b) Using the self-consistency property (4.1) avoids large growth in the Jacobians, enabling successful learning of the correct dynamics.
(c) Similar performance in 100dd Müller- Brown example suggests increased dimensionality does not degrade performance when intrinsic difficulty remains unchanged.
Figure 7: Comparison of the trajectories obtained using our method, in the Müller-Brown examples.
100dd Müller-Brown Example
To assess performance in higher dimensions, we also report results for an augmented Müller-Brown potential in 100 dimensions (following a similar 5dd example in Sipka et al. (2023)). The first two dimensions match the standard Müller-Brown potential above, and the added dimensions follow a quadratic potential. We use a larger network with hidden dimensions [256,256,256] and time-embedding hidden dimensions [128,128], and train each method for 8000 steps; otherwise, we use the same hyperparameters as above.
For the obtained paths, in Table 4 we report the same quantitative results as in Table 3. To make comparable with those in Table 3, we use the first two components of the obtained trajectories to evaluate the metrics, so the additional dimensions only influence training and not evaluation (otherwise the KL computations are dominated by the behaviour in the added dimensions). Trajectories for the first two dimensions using our method are visualised in Figure 7(c).
We find that our method displays largely similar performance in this higher-dimensional example compared to the standard 2dd case, and the ability to learn the bridge does not appear to deteriorate significantly. This suggests the difficulty of the problem is primarily governed by the complexity of the transition path rather than the ambient dimensionality, and that our approach appears not to be substantially impacted by increases in dimensionality when the underlying problem structure remains unchanged.